The decision variables represent the amounts of ingredients 1, 2, and 3 to put into a blend. The objective function represents profit. The first three constraints measure the usage and availability of resources A, B, and C. The fourth constraint is a minimum requirement for ingredient 3.

 

LP Problem

 

Maximize 4x1 + 6x2 + 7x3


s.t.

3x1+2x2+5x3 ≤ 120

  x1+3x2+3x≤  80

5x1+5x2+8x3 ≤ 160

    x3 ≥  10

         x1, x2 , x≥ 0

 

Answer the following questions.

a. How much of ingredient 1 will be put into the blend?

b. How much of ingredient 2 will be put into the blend?

c. How much of ingredient 3 will be put into the blend?

d. How much resource A is used?

e. How much resource B will be left unused?

f. What will the profit be?

g. What will happen to the solution if the profit from ingredient 2 drops to 4?

h. What will happen to the solution if the profit from ingredient 3 increases by 1?

i. What will happen to the solution if the amount of resource C increases by 2?

j. What will happen to the solution if the minimum requirement for ingredient 3 increases to 15?


1
Expert's answer
2021-03-12T05:42:55-0500

Using Simplex method online calculator (https://cbom.atozmath.com/CBOM/Simplex.aspx?q=sm), we get:

a) "x_1=0"

b) "x_2=16"

c) "x_3=10"

d) "2\cdot16+3\cdot10=62"

e) "80-(3\cdot16+3\cdot10)=2"

f) Profit: "6\cdot16+7\cdot10=166"


g) There is change in the objective function:

"4x_1+4x_2+7x_3"

Then:

"x_1=x_2=0,x_3=20,z=140"


h) Change in the objective function:

"4x_1+4x_2+8x_3"

Result:

"x_1=0,x_2=16,x_3=10,z=176"


i) Change in the 3rd constraint:

"5x_1+5x_2+8x_3 \u2264 162"

Result:

"x_1=0,x_2=16.4,x_3=10,z=168.4"


j) Change in the 4rd constraint:

"x_3 \u2265 15"

Then:

"x_1=0,x_2=8,x_3=15,z=153"